For the gaseous reaction, N2 + 3H2 ↔ 2NH3  the partial pressures of H2 and N2 are 0.4 and 0.8 atm respectively. The total pressure of the entire system is 2.8 atm. What will be the value of Kp if all the concentrations are given in atmospheres??

Grade:12

4 Answers

Sudheesh Singanamalla
13 years ago

N2 + 3H2 ↔ 2NH3

partial pressures of H2 and N2 are 0.4 and 0.8 atm

so partial pressure of 2NH3  = 1.6 atm

Kp= [NH3] ^2 / ( [N2] [ H2]^3 )

= 1.6 ^ 2 / 0.8 * 0.4 ^3

= 2.56 / 0.0512

= 50

Please approve !


ajinkya bhole
13 years ago

if partial pressures tht u hav givn fr h2 nd n2 r at equilibrium thn...

presure of nh3=1.6

so kp=(1.6)^2/0.8*(0.4)^3

vikas askiitian expert
13 years ago

N2      +      3H2     =>    2NH3

0.4            0.8                  P                 (at equilibrium)

sum of all pressure = P(total) = 1.2 + p = PT

  PT  = 2.8 (given)  so

 1.2 + P = 2.8

         pNH3 = 1.6

now Kp =  ( pNH3)2 /(PN2)(pH2)3

           = (1.6)2/(0.4)(0.8)3

          Kp =  12.5 (atm)-2

Aiswarya Ram Gupta
13 years ago

thnx to al......